OpenStudy (anonymous):

PLEASE HELP!! I dont understand!! Think about a kayaking race, a round trip airplane flight, a shopper on an escalator, or a swimming competition. Choose an activity where a rational expression can represent the total or difference of time. For example, what is the time to go round trip on a hot air balloon ride? Use full sentences to describe the real world situation. Write the addition or subtraction rational expression to represent your situation. Simplify the expression, showing all the work or explaining the steps using complete sentences

OpenStudy (amistre64):

unit rate * time = distance traveled that might be useful, not sure how to esssay up the results tho

OpenStudy (amistre64):

total time = time to + time return hotair time = distance/speed of ballon + distance/speed of car(walk,boat,helicopter) ride back

OpenStudy (anonymous):

well lets say it was a 8 hour roundtrip from virginia to california. what would i do next?

OpenStudy (amistre64):

what is the distance travled from V to C ?

OpenStudy (anonymous):

let me see

OpenStudy (amistre64):

lets avoid "round trip" time for now; lets split it up into time to, and time from

OpenStudy (anonymous):

2230 miles

OpenStudy (amistre64):

hmm, to make these into expressions; lets say it is some distance greater then 10 :) distance = 10+x

OpenStudy (amistre64):

how are we gettting from V to C? plane, car, train, amc gremlin?

OpenStudy (anonymous):

plane

OpenStudy (amistre64):

does it take 4 hours to get from V to C in a plane? just to make this thing as realiztic as possible

OpenStudy (anonymous):

ill make sure

OpenStudy (anonymous):

2 and a half hours

OpenStudy (amistre64):

\[total_{time}=\frac{x+10}{2230/2.5}+time_{from}\] \[total_{time}=\frac{x+10}{892}+time_{from}\]

OpenStudy (amistre64):

if the same method of travel coming back is used we just double the it

OpenStudy (amistre64):

\[total_{time}=\frac{2(x+10)}{892}\]5 hours is the total time to go round trip in this case

OpenStudy (amistre64):

im not sure why they want you to apply rational expressions ....

OpenStudy (anonymous):

okay so what is he rational expression i use?

OpenStudy (anonymous):

the

OpenStudy (amistre64):

lets say x represents a distance between V and C that is 10 miles away from the starting point

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

\[total_{time}=\frac{2(x-10)}{892}\]

OpenStudy (amistre64):

a rational expression is a glorified fraction

OpenStudy (anonymous):

okay, so then all i have to do is solve that?

OpenStudy (amistre64):

still not sure HOW they want you to write it out tho .... the total time would be the solution to that yes; but im not sure if that is a correct rendering of the situation :/

OpenStudy (anonymous):

well it says i have to add or subtract in the rational expression?

OpenStudy (amistre64):

im not sure if that is read "within" the rational expression; or, if its the addition/subtraction OF rational expressions ....

OpenStudy (amistre64):

if the total time taken to make the circuit is dependant upon different modes of transportation .. that would give us an addition of rationals

OpenStudy (anonymous):

im not sure which one they want..

OpenStudy (amistre64):

me either ....

OpenStudy (amistre64):

see if jchavez has any input :) since they said they are working on the same material

OpenStudy (anonymous):

okay, i'll use the 2(x-10)/892 & okay i'll see(: